OpenStudy (anonymous):

i dont understand this. evaluate the trig functions using its period as an aid. sin5π what is this "period"

OpenStudy (lgbasallote):

i believe period is one revolution which in radians is \(2\pi\)

OpenStudy (lgbasallote):

so \(5 \pi\) is 2 revolutions and a half

Parth (parthkohli):

If you want to be explained in degrees. 360 + 360 + 180 720 + 180 720 degrees = 0 180 = pi degrees.

OpenStudy (lgbasallote):

so \(\sin (5\pi) = \sin (\pi)\) make sense?

Parth (parthkohli):

Just calculate: \(\sin\pi\) \(\cos \pi\) \(\tan \pi\)

Parth (parthkohli):

Or in degrees: \(\sin 180^{\circ}\) \(\cos 180^{\circ}\) \(\tan 180^{\circ}\)

OpenStudy (lgbasallote):